If x is real show that {1/(x+1)} + {1/(3x+1)} - {1/(x+1)(3x+2)} can have all numerical values except those lying between 1 and 4.
 
Please give the solution of above problem.
 
Warm Regards
Nafis Akhtar

Grade:11

1 Answers

Susmita
5 years ago
\frac{1}{x+1}+\frac{1}{3x+1}-\frac{1}{(x+1)(3x+2)}
 
Now see if x=-1 or x=-2/3 or x=-1/3 then the denominator of the corresponding terms becomes zero and the expression becomes infinity.
So these 3 values of x is not allowed.I do not understand why it has to be between 1 to 4.